OpenStudy (anonymous):

A square patio has an area of 121 square yards. What is the length of this patio? ______ yards

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Any more information?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thats all it gave me

OpenStudy (igreen):

The formula for the area of a square is: A = a^2 Plug in what we know: 121 = a^2 Find the square root of 121

OpenStudy (igreen):

That will be your answer.

OpenStudy (anonymous):

i dnt speek math i kinda suck at it

OpenStudy (igreen):

\(\sqrt{121} = Answer\)

OpenStudy (anonymous):

better lol

OpenStudy (igreen):

xD

jaynator495 (jaynator495):

xD

OpenStudy (anonymous):

hlp

OpenStudy (anonymous):

no-one

OpenStudy (anonymous):

all just sitting here to sit here